A bottle full of water has mass of 45g when full of mercury has mass of 360g when empty mass is 20g .calculate mass of mercury,water.relative density of mercury.density of mercury if density of water is 1000kgm-3


1
Expert's answer
2021-09-17T15:32:50-0400

The mass of mercury is


"m_m=360-20=340\text{ g}."


The mass of water is


"m_w=45-20=25\text{ g}."

Since the volume of water and mercury is the same, we have


"\rho_m=\frac{m_m}{V},\\\space\\
\rho_w=\frac{m_w}{V},\\\space\\
r=\frac{\rho_m}{\rho_w}=\frac{m_m}{m_w}=13.6."

The density of mercury is


"\rho_m=r\rho_w=13.6·1000=13600\text{ kg/m}^3."